# Underactuated Control of Multiple Soft Pneumatic Actuators via Stable Inversion

### v1

@@ III-C Controller Design and Analysis
[14] Consider the block diagram in Fig. 3 with (22). Then the bounded $U_{fb}$ yields $\parallel y_{d}(t)-\tilde{y}(t)\parallel_{\infty}\rightarrow 0$ $\textit{iff}~{}\tilde{Y}(s)\in\textnormal{Im}_{\mathbb{R}(s)}(P)$.
Figure 3: The block diagram of the algebraic controller including the feedforward control and feedback loop.
For simplicity, we take ${\Delta}{W_{T}}$ as $\Delta$, and consider $\tilde{P}(s)\in\Pi$, $Y_{d}=PU_{ff}$, and $U_{c}(s):=U_{ff}-U_{fb}$ (by referencing Fig. 3)
$\tilde{P}(s)U_{c}(s)=\tilde{Y}(s)=PU_{c}+\Delta_{u}PU_{c}$ (24)
$PU_{ff}+\Delta_{u}PU_{ff}-PU_{fb}-\Delta_{u}PU_{fb}=\tilde{Y}(s)$ (25)
$\Rightarrow PU_{fb}=\Delta_{u}PU_{ff}-\Delta_{u}PU_{fb}$ (26)
The feedback loop will compensate for the model errors. The detail proof of Theorem 2 and 3 can be referenced in [14].
The model inversion and feedback loop are introduced to achieve accurate tracking of the system. The next problem is whether the system performance can be achieved by underactuated control as depicted in (5). That is, we control multiple soft actuators with a single input pressure. Here, we intend to synchronize the motions of the multi-finger gripper to reach stable grasping [7]. The desired response of each finger is set as the same so the (20) becomes
$\begin{bmatrix}Y_{d}(s)\\ \vdots\\ Y_{d}(s)\end{bmatrix}$ $=\begin{bmatrix}P_{1}(s)\\ \vdots\\ P_{n}(s)\end{bmatrix}{U(s)}$ (27)
$\Rightarrow{Y}_{d}(s)$ $=P(s)U(s)$ (28)
where ${Y}_{d}(s)\in\textnormal{Im}_{\mathbb{R}(s)}(P)$. We are going to show that this problem is solvable and the solution exists by mathematics inference. Since the $P(s)$ has full rank = 1, the $P(s)$ is invertible and exists left inverse matrix according to the Theorem 2. The inverse matrix is $P\textsuperscript{\textdagger}(s)=(P^{*}(s)P(s))^{-1}P^{*}(s)$. The controller is obtained by $U(s)=H(s)P\textsuperscript{\textdagger}(s){Y}_{d}(s)$. Therefore, there exists input $U(s)$ that makes the ${Y}_{d}(s)$ achievable.
Furthermore, if the system is controllable, the controllability matrix of the (5) is defined as $M_{c}=[A~{}BA~{}BA^{2}~{}\cdots]$. The set $\mathbb{S}_{n}\subset\mathbb{R}$ is assumed as the synchronization subspace [31]. Now we define the controllability of the (28).
Definition 2.
The system matrix $P(s)$ in (28) is minimal and controllable if $\mathbb{S}_{n}\subset M_{c}$.
The single input may fail to synchronize the motions of soft fingers with the same dimensional parameters because of the uncertainty of the soft materials. The deformation curves of soft materials exhibit high uncertainty when they have a slow deformation rate [20]. In contrast, the curves demonstrate much less uncertainty when their deformation rate is high. Similarly, when the higher pressure changing rate is applied to soft fingers, they show less uncertainty, align with nominal behaviors, and tend to have synchronized motions. This property influences the controllability of the synchronization of the multi-finger soft gripper (28) and the existence of the subspace $\mathbb{S}_{n}$.
Figure 4: Several open-loop responses of both soft pneumatic actuators (soft fingers) are demonstrated in (a). The robustness weight selection based on the modeling errors can be seen in (b).
@@ IV Experimental Evaluation
In previous sections, we introduce the mechatronic system, specifically the multi-finger soft gripper, alongside system dynamical models and control algorithms. Prior to experimentation, preliminary tests are conducted using analytical software to assess the feasibility of the algebraic controller. Subsequently, a series of experiments are executed to evaluate the practicality of the proposed control approach. Additional disturbance tests are then performed to evaluate the robustness of the controller.
@@ IV-A Preliminary Evaluations
Some definitions and assumptions are applied in Sec. II and III, so this subsection intends to evaluate that the definitions and assumptions are valid before the simulations and experimentation.
@@ IV-A1 Model Evaluation
The analytical model matrix is built for the mechatronic system as the $P(s)$ in (29). Based on the discussion in Sec. III-C, the model has uncertainty and will cause modeling errors due to the uncertain soft materials. The bounded model uncertainty is assumed, $\parallel\Delta\parallel_{\infty}\leq 1$ in (22). Several step responses of both soft fingers are conducted to evaluate whether the model uncertainty is bounded.
Figure 4 (a) demonstrates the repeated step responses of both soft actuators, the first (blue) and second (green) element of the $P(s)$ in (29). The system performance varies due to the properties of soft materials [7]. The singular value of the bounded constraint can be found in Fig. 4 (b). The system perturbation of two soft actuators is bounded. Specifically, it is observed that $\parallel C_{n}-C_{nominal}\parallel~{}\leq 14.3\%$ and $\parallel K_{n}-K_{nominal}\parallel~{}\leq 5.9\%$ by system identification of the step responses in Fig. 4 (a). Two soft actuators, both elements of the $P(s)$ in (29), show a similar result. We can conclude that the perturbations of the two systems are bounded.
@@ IV-A2 Controllability Evaluations
Another preliminary evaluation is needed before the experimentation. The system equation of (4) should be controllable and observable. The controllability matrix $M_{c}=[A~{}BA~{}BA^{2}~{}\cdots]$ and the observability matrix $M_{o}=[C~{}CA~{}CA^{2}~{}\ldots]^{T}$ of (4) both have full rank, so the full system is controllable and observable. Since the full system is controllable and observable, the system realization is also minimal. This evaluation matches the Assumption 1. The designed algebraic controller is valid.
Figure 5: The simulation results of the two-finger gripper are shown in (a). The sensor noise and disturbance are considered and displayed in (b) and (c).
@@ IV-B Simulation Results
Section II-B introduces the general case of the problem (5). The system matrix $P(s)$ comprises $n$ elements. Simulations will be conducted on multi-finger soft gripper systems to assess different scenarios, with $n$ taking values of 2 and 3 for $P(s)$, respectively. The simulations help us understand if the controller can work and make adjustments. For example, the low pass filter $H(s)$ of Theorem 4 will be adjusted to achieve better tracking performance.
IV-B1 Two-finger Gripper Simulation
@@
The first simulation is performed on a two-finger gripper, which includes two soft actuators (soft fingers) and a single syringe pump. The desired output is selected as a step response $\frac{\pi/3}{s}$ $\in\textnormal{Im}_{\mathbb{R}(s)}(P)$ considering the applications. The desired output $(\pi/3)$ may lead to the collision of the two soft fingers as Fig. 1, but we aim to check the performance of the control algorithm. The system equation is written as
$\begin{bmatrix}\frac{\pi/3}{s}\\ \frac{\pi/3}{s}\end{bmatrix}$ $=\begin{bmatrix}\frac{7.831}{s^{3}+2.66s^{2}+3.61s}\\ \frac{7.831}{s^{3}+2.45s^{2}+3.06s}\end{bmatrix}{U(s)}$ (29)
The controller is obtained based on the Theorem 2 and 3, so the $H(s)$ is designed as $5/s$ and $P\textsuperscript{\textdagger}(s)=[\frac{s^{3}+2.66s^{2}+3.61s}{15.66},~{}\frac{s^{3}+2.45s^{2}+3.06s}{15.66}]$.
@@
This controller is able to synchronize the motions of the two fingers within the soft gripper as in Fig. 5 (a). The settling time of both fingers is below 1 $sec$. To further evaluate the performance of the controller, sensor noise and disturbance are added as Fig. 5 (b) and (c). The feedback loop can compensate for the sensor noise as Fig. 5 (b) The root-mean-square error (RMSE) of the sensor noise is around $\pm$ 2 $deg$ [7]. The RMSE of the feedforward control only and feedforward and feedback control of simulations are approximately 5 deg and 2.3 $deg$ respectively. Thus, the feedback loop outperforms in handling the noise.
@@
The disturbance here is regarded as the soft fingers are hitting by an external force. The results show that the controller is capable of adjusting the system back to the reference faster than feedforward control as Fig. 5 (c). Moreover, the amplitude of the errors of feedforward controller $(0.4~{}rad)$ is larger than that of the feedback loop $(0.2~{}rad)$. The simulation results validate the Theorem 3 and (26). The experimental results can be referenced in Sec. IV-E and IV-F.
Figure 6: The simulation result of the three-finger soft gripper is demonstrated. Their models are different but their motions are synchronized by the algebraic controller.
IV-B2 Three-finger Gripper Simulation
The algebraic controller displays the desired performance in the two-finger gripper simulations. Next, the simulation is extended to a three-finger gripper. The reference is the same and it is $\frac{\pi/3}{s}$ $\in\textnormal{Im}_{\mathbb{R}(s)}(P)$. This desired output will also lead to collision of the three fingers but the simulation here aims to evaluate the performance of the controller.
$\begin{bmatrix}\frac{\pi/3}{s}\\ \frac{\pi/3}{s}\\ \frac{\pi/3}{s}\end{bmatrix}=\begin{bmatrix}\frac{7.831}{s^{3}+2.66s^{2}+3.61s}\\ \frac{7.831}{s^{3}+2.76s^{2}+3.88s}\\ \frac{7.831}{s^{3}+2.45s^{2}+3.06s}\end{bmatrix}{U(s)}$ (30)
The $H(s)$ is designed as $5/s$ and $P\textsuperscript{\textdagger}(s)=[\frac{s^{3}+2.66s^{2}+3.61s}{23.49},~{}\frac{s^{3}+2.76s^{2}+3.88s}{23.49},~{}\frac{s^{3}+2.45s^{2}+3.06s}{23.49}]$.
The simulation results are demonstrated in Fig. 6. The controller is able to synchronize the three-finger soft gripper and the motions of the three fingers are aligned. Their response time is all below 1 $sec$.
IV-B3 Summary
@@
The simulation results on two-finger and three-finger soft grippers endorse the feasibility of the algebraic controller. The motions of fingers are synchronized in both the transient and steady states. The synchronization will enable the grasping success rate when the soft gripper is applied to manipulate various objects. The controller will be applied to the real soft gripper to evaluate the applicability of the algebraic control algorithms.
@@ IV-C Experimental Setup
Figure 1 illustrates both the experimental arrangement and the signal flow diagram. The two-finger soft gripper is used to conduct the experiments with a single syringe pump. The soft actuators are fabricated by molds as illustrated in Fig. 2 (b). The motions of the soft fingers are driven by the syringe pump [24], which is actuated by a stepper motor. An air pressure sensor (Walfront, Lewes, DE) with a sensing range of 0 to 80 $psi$ is utilized to detect the air pressure for open-loop control. Additionally, each soft finger contains a flex sensor (Walfront, Lewes, DE) inside to monitor its bending angle, facilitating feedback control. The flex sensor is a resistive type sensor and has a sensing range of 100 $deg$ and sensing error is approximately 2 deg (root-mean-square error). Both sensors are synchronized with Arduino MEGA 2560 (SparkFun Electronics, Niwot, CO), which is based on the Microchip ATmega 2560. The controller algorithms are programmed in MATLAB/Simulink which is communicated with the Arduino MEGA 2560 to process feedback signals and generate control commands for the mechatronic system. The model and controller are discretized in the analytical software with a sampling time of 100 $ms$. The ”Real System” block in Fig. 3 is replaced by the real soft gripper.
Figure 7: The visualization of the open-loop test and LQR control of the two-finger gripper is asynchronized as displayed in (a). The algebraic control result is demonstrated in (b). The motions of the two fingers are synchronized compared to the results in (a). The disturbance test is disclosed in (c) and the controller can handle the external disturbance.
@@ IV-D Open-loop & Closed-loop Tests
Prior to implementing the algebraic controller, the open-loop tests attempt to visualize the open-loop responses of the soft fingers. The soft fingers have the same dimensional parameters such as height, width, length, etc. Nonetheless, their system parameters of (19) are different which leads to asynchronized motions. In applications, the open-loop control results in grasping failure of the soft gripper [7] due to the asynchronized motions of its fingers. The open-loop test results are illustrated in Fig. 7 (a).
@@
The blue dashed line represents the response of the left finger (SPA 1) in Fig. 1 while the green dashed line denotes the right finger (SPA 2) in Fig. 1. The left finger is active and responds faster. By contrast, the right finger has relatively slow responses. Their steady states are also different due to different system parameters of (19). The left finger reaches approximately 47 deg while the right finger reaches around 40 deg with a reference 45 deg.
Moreover, the closed-loop test is performed by using a single LQR controller for two soft fingers. The accuracy and response time of both fingers are improved; however, their motions are not synchronized in Fig. 7 (a). The right finger has a steady-state error of around 4–5 deg compared to the reference. Thus, a single LQR is unable to synchronize the motions of two fingers due to their model variations.
IV-E Stable Inversion Control Tests

### v2

@@ III-C Controller Design and Analysis
[21] Consider the block diagram in Fig. 3 with (22). Then the bounded $U_{fb}$ yields $\parallel y_{d}(t)-\tilde{y}(t)\parallel_{\infty}\rightarrow 0$ $\textit{iff}~{}\tilde{Y}(s)\in\textnormal{Im}_{\mathbb{R}(s)}(P)$, where $y_{d}(t)$ has the same definition in Theorem 2 and $\tilde{y}$ is the real system response with feedback loop.
For notational simplicity, we take ${\Delta}{W_{T}}$ in (22) as $\Delta_{P}$, and consider $\tilde{P}(s)\in\Pi$, $Y_{d}=PU_{ff}$, and $U_{c}(s):=U_{ff}-U_{fb}$ (by referencing Fig. 3)
$\tilde{P}(s)U_{c}(s)=\tilde{Y}(s)=PU_{c}+\Delta_{P}PU_{c}$ (26)
$PU_{ff}+\Delta_{P}PU_{ff}-PU_{fb}-\Delta_{P}PU_{fb}=\tilde{Y}(s)$ (27)
$\Rightarrow PU_{fb}=\Delta_{P}PU_{ff}-\Delta_{P}PU_{fb}$ (28)
The feedback loop and $U_{fb}$ will compensate for the model errors ($\Delta_{P}$). The proof of Theorem 2 and 3 are given in [21].
The model inversion and feedback loop are introduced to achieve accurate tracking of the system. The next problem is whether the system performance can be achieved by underactuated control as depicted in (24). That is, multiple soft actuators are controlled with a single input pressure. Here, the goal is to control the motions of the multi-finger gripper to reach stable grasping [7]. The desired response of each finger is assigned as $Y_{d}$ in (24) and the ${Y}_{d}(s)\in\textnormal{Im}_{\mathbb{R}(s)}(P)$. This problem will be shown to be solvable and the solution to exist through mathematical inference. Since the $P(s)$ has full rank = 1, the $P(s)$ is invertible and exists left inverse matrix according to the Theorem 2. The pseudo-inverse is $P\textsuperscript{\textdagger}(s)=(P^{T}(s)P(s))^{-1}P^{T}(s)$. The controller is obtained by $U^{a}(s)=H(s)P\textsuperscript{\textdagger}(s){Y}_{d}(s)$. Therefore, there exists input $U(s)$ that makes the ${Y}_{d}(s)$ achievable.
Figure 4: Several open-loop responses of both soft pneumatic actuators (soft fingers) are demonstrated in (a). The robustness weight selection of both soft fingers based on the modeling errors can be seen in (b).
@@ IV Experimental Evaluation
In previous sections, we introduce the mechatronic system, specifically the multi-finger soft gripper, alongside system dynamical models and control algorithms. Prior to experimentation, preliminary tests are conducted using MATLAB®/Simulink to assess the feasibility of the stable inversion algorithm. Subsequently, a series of experiments are executed to evaluate the practicality of the proposed control approach. An additional disturbance test is then performed to evaluate the robustness of the controller.
@@ IV-A Preliminary Evaluations
Some definitions and assumptions are applied in Sec. II and III. This subsection intends to evaluate that the definitions and assumptions are valid before the simulations and experimentation.
@@ IV-A1 Model Evaluation
The analytical model matrix is built for the mechatronic system as $P(s)$ in (31). Based on the discussion in Sec. III-C, the model has uncertainty and will cause modeling errors due to the uncertain soft materials. The bounded model uncertainty is assumed, $\parallel\Delta\parallel_{\infty}\leq 1$ in (22). Several step responses of both soft fingers are conducted to evaluate whether the model uncertainty is bounded.
Figure 5: The step responses of different speeds of SPA1 (left finger in Fig. 1) are shown in (a). The robustness weight selection of SPA1 with high and slow speeds based on the modeling errors can be seen in (b).
Figure 4 (a) demonstrates the repeated step responses of the two soft actuators, the first (blue) and second (green) element of the $P(s)$ in (31). The system performance varies due to the properties of soft materials [7]. The singular values of the bounded constraint in (22) can be found in Fig. 4 (b). The system perturbation of two soft actuators is bounded. Specifically, based on the (16), it is observed that $\parallel C_{n}-C_{nominal}\parallel~{}\leq 14.3\%$ and $\parallel K_{n}-K_{nominal}\parallel~{}\leq 5.9\%$ by performing system identification of the step responses in Fig. 4 (a). The $C_{nominal}$ and $K_{nominal}$ are the average of the identified model of those multiple step responses. Two soft actuators, both elements of the $P(s)$ in (31), show a similar result. It is concluded that the perturbations of the two systems are bounded, confirming the validity of the model evaluation. This evaluation holds within the excited frequency range from 6 to 63 $rad/s$ as the yellow region in Fig. 4 (b), determined by performing the Fast Fourier transform on the input reference as shown in Fig. 4 (a).
@@ IV-A2 Controllability Evaluations
Another preliminary evaluation is needed before the experimentation. The system equation of (6) should be controllable and observable to ensure that the system realization is minimal. Note that if the two fingers are identical, the realization is uncontrollable. In this case, the motion of the two fingers is always synchronized. If the controllability matrix $M_{c}=[A~{}BA~{}BA^{2}~{}\cdots]$ and the observability matrix $M_{o}=[C~{}CA~{}CA^{2}~{}\ldots]^{T}$ of (6) both have full rank, the full system is controllable and observable. Since the full system is controllable and observable, the system realization is also minimal. This evaluation matches the Assumption 1. a. The proposed controller design is valid.
Figure 6: The simulation results of the under-actuated control of the two-finger gripper are depicted in (a). The sensor noise and external disturbance are considered and visualized in (b) and (c).
IV-A3 Uncertainty Band of Soft Finger
It is observed that the soft materials exhibit larger uncertainty as discussed in Remark 1 of Sec. III-C. The experimental evaluation of this property is conducted and visualized in Fig. 5 (a) and (b). The left finger in Fig. 1 is used to conduct this evaluation. Two input commands with different speeds are applied to do step response tests of a soft finger. The slower-speed input command excites a larger uncertainty band (red lines). The standard deviation of the steady-state error is 2.41 $deg$, while that of higher-speed input command is 1.14 $deg$ (blue lines). The input references in Fig. 5 (a) are converted to the frequency domain using the Fast Fourier transform to validate the excited frequency region, as the yellow region in Fig. 5 (b). This soft actuator’s frequency region spans from approximately 6 to 50 $rad/s$. Figure 5 (b) illustrates that the magnitude variations of the soft actuator converge at higher frequencies, approaching the nominal dynamics. These results confirm that the soft actuator demonstrates reduced uncertainty and approximates nominal dynamics at higher speeds (or frequencies), consistent with the observations in the time domain. This property plays a vital role in the multiple soft finger underactuated control since soft fingers have different dynamical models. Higher speeds may reduce the uncertainty band and help coordinate multiple soft fingers within a soft gripper.
@@ IV-B Simulation Results
Section II-B introduces the general case of the problem (6). The system matrix $P(s)$ comprises $n_{y}$ elements. Simulations will be conducted on multi-finger soft gripper systems to assess different scenarios, with $n_{y}$ taking values of 2 for $P(s)$. The simulations help us understand if the controller can work and regulate the system output. Besides, the low pass filter $H(s)$ of Theorem 2 will be adjusted to achieve better tracking performance.
@@
The simulation is performed on a two-finger gripper, which includes two soft actuators (soft fingers) and a single syringe pump. The ideal desired output is selected as a step response ${Y}_{d}=[\frac{0.333\pi}{s},\frac{0.333\pi}{s}]^{T}$. As the elements in $P(s)$ are nonidentical in (31), the ${Y}_{d}$ is not in the image space of $P$ and it can be factorized as [21]
$Y_{d}=\underset{\text{Im}(P)}{\textnormal{proj}}(Y_{d})+\textnormal{res}(Y_{d})$ (29)
$\underset{\text{Im}(P)}{\textnormal{proj}}(Y_{d})=\sum_{n=1}^{r}\langle{Y_{d}},{q_{i}}\rangle{q_{i}}$ (30)
where ${q_{i}}s$ is defined in (2). The $\textnormal{res}(Y_{d})$ can be obtained by $Y_{d}-\textnormal{proj}(Y_{d})$ based on (29) and (30). Note that this obtained desired output vector may lead to the collision of the two soft fingers as Fig. 1, but we aim to check the performance of the control algorithm. With the $\textnormal{res}(Y_{d})$, $\textnormal{proj}(Y_{d})=[0.303\pi/s,0.357\pi/s]^{T}$ is in the range space of $P$. The system equation is written as
$\begin{bmatrix}\frac{0.303\pi}{s}\\ \frac{0.357\pi}{s}\end{bmatrix}$ $=\begin{bmatrix}\frac{7.831}{s^{3}+2.66s^{2}+3.61s}\\ \frac{7.831}{s^{3}+2.45s^{2}+3.06s}\end{bmatrix}{U(s)}$ (31)
The systems in (31) will have steady-state tracking errors, $0.030\pi$ (SPA 1) and $-0.024\pi$ $rad$ (SPA 2) respectively compared to the desired reference ($\pi/3~{}rad$). The controller is obtained based on the Theorem 2 and 3, so the $P\textsuperscript{\textdagger}(s)=[\frac{s^{3}+2.66s^{2}+3.61s}{15.66},~{}\frac{s^{3}+2.45s^{2}+3.06s}{15.66}]$ and the $H(s)$ is designed by loop-shaping the element of P(s) with desired equation $4.8/s$, and $H(s)=\frac{1009s^{3}+2.52e5s^{2}+6.829e5s+1.027e6}{s^{6}+282.7s^{5}+6.14e3s^{4}+8.022e4s^{3}+4.335e5s^{2}+8.948e5s+1.027e6}.$
@@
This controller is able to coordinate the motions of the two fingers within the soft gripper as Fig. 6 (a). The settling time of each finger is approximately 0.6 $sec$, making the soft gripper comparable to traditional grippers. The tracking error of the two fingers is within 1 $deg$ compared to the $\textnormal{proj}(Y_{d})$. To further evaluate the performance of the controller, sensor noise and disturbance are added as Fig. 6 (b) and (c). The feedback loop can compensate for the sensor noise as Fig. 6 (b). The root-mean-square error (RMSE) of the sensor noise is set as $\pm$ 2 $deg$ [7]. The RMSE of the two fingers is within 0.5 $deg$, so the sensor noise does not affect the systems’ output.
@@
The disturbance here is regarded as the soft fingers are hitting by an external force. The results show that the controller is capable of adjusting the system back to the reference as Fig. 6 (c). The amplitude of the errors for both fingers is within 0.1 $rad$, so the feedback control can handle the disturbance. The simulation results validate the Theorem 3 and (28). The experimental results can be referenced in Sec. IV-E and IV-F.
@@
The simulation results on two-finger soft grippers endorse the feasibility of the proposed controller algorithm. The motions of fingers are coordinated. These dynamics will be beneficial to grasping tasks when the soft gripper is applied to manipulate various objects. The controller will be applied to the real soft gripper to evaluate the applicability of this control algorithm.
@@ IV-C Experimental Setup
Figure 1 illustrates both the experimental arrangement and the signal flow diagram. The two-finger soft gripper is used to conduct the experiments with a single syringe pump. The soft actuators are fabricated by molds as illustrated in Fig. 2 (b). The motions of the soft fingers are driven by the syringe pump as Fig. 1 [28], which is actuated by a stepper motor. A DM320T digital stepper driver (StepperOnline, New York, NY) is utilized to trigger the stepper motor. An air pressure sensor (Walfront, Lewes, DE) with a sensing range of 0 to 80 $psi$ is utilized to detect the air pressure for open-loop control. Additionally, each soft finger contains a flex sensor (Walfront, Lewes, DE) inside to monitor its bending angle, facilitating feedback control. The flex sensor is a resistive type sensor and has a sensing range of 100 $deg$ and a sensing error of approximately 2 $deg$ (root-mean-square error). Both sensors are synchronized with Arduino MEGA 2560 (SparkFun Electronics, Niwot, CO), which is based on the Microchip ATmega 2560. The controller algorithms are programmed in MATLAB®/Simulink which is communicated with the Arduino MEGA 2560 to process feedback signals and generate control commands for the mechatronic system. The model and controller are discretized in the analytical software with a sampling time of 100 $ms$. The “Real System” block in Fig. 3 is replaced by the real soft gripper.
Figure 7: The visualization of the open-loop test of the two-finger gripper is displayed in (a). The responses of using the proposed controllers designed for different bandwidths are demonstrated in (b). The motions of the two fingers are coordinated compared to the results in (a). The disturbance test is depicted in (c) and the controller can handle the external disturbance.
@@ IV-D Open-loop & Closed-loop Tests
Prior to implementing the proposed controller, the open-loop test attempts to visualize the open-loop responses of the soft fingers. The soft fingers share the same dimensional parameters such as height, width, and length. However, their system parameters in equation (21) differ, resulting in distinct motions. In applications, the open-loop control results in inconsistent motions of multiple fingers which is not beneficial for manipulation tasks. The open-loop test results are illustrated in Fig. 7 (a).
@@
The desired reference is set as 45 $deg$ ($\pi$/4 $rad$) to avoid real collision between two fingers. Based on the (29) and (30), the desired references for left and right fingers are 48.24 and 40.86 $deg$, respectively. The blue dashed line represents the response of the left finger (SPA 1) in Fig. 1 while the green dashed line denotes the right finger (SPA 2) in Fig. 1. The left finger is active and responds faster. By contrast, the right finger has relatively slow responses. Their steady states are also different due to different system parameters of (31). The left finger reaches approximately 50 $deg$ while the right finger reaches around 39.5 $deg$.
IV-E Underactuated Control Tests
The stable inversion algorithm in Sec. III-C is implemented to control the two fingers in this subsection. The system equation is similar to (31) in Sec. IV-B and is described as
$\begin{bmatrix}\frac{0.227\pi}{s}\\ \frac{0.268\pi}{s}\end{bmatrix}$ $=\begin{bmatrix}\frac{7.831}{s^{3}+2.66s^{2}+3.61s}\\ \frac{7.831}{s^{3}+2.45s^{2}+3.06s}\end{bmatrix}{U(s)}$ (32)
where $P\textsuperscript{\textdagger}(s)$ and $H(s)$ are the same as shown in Sec. IV-C, and the desired reference is set as 45 $deg$ ($\pi$/4 $rad$). According to the (29) and (30), the $Y_{d}=[0.25\pi/s,0.25\pi/s]^{T}$ is factorized as $\textnormal{proj}(Y_{d})$ and $\textnormal{res}(Y_{d})$. The $\textnormal{proj}(Y_{d})=[0.227\pi/s,0.268\pi/s]^{T}$ is in the range space of $P(s)$, and the steady-state tracking errors of systems in (32) are $0.023\pi$ and $-0.018\pi$ $rad$ respectively compared to the desired reference $\pi/4~{}rad$.
The result is demonstrated in Fig. 7 (b). Two soft fingers reach their steady states at nearly the same time, and their tracking errors are below 1 $deg$ compared to the $\textnormal{proj}(Y_{d})$. The settling time is around 0.7 $sec$, which is better than our previous research by using an optimal controller [7]. The response time enables the soft gripper to be comparable to rigid grippers. Besides, their transient states are nearly synchronized with around 2 $deg$ differences, which support the evaluation of Sec. IV-A3.
Additionally, the bandwidth of $H(s)$ affects the systems’ responses, so another controller is designed whose $H(s)$ has narrower bandwidth. The $H(s)$ is designed by loop-shaping the element of P(s) with desired equation $2/s$, and $H(s)=\frac{49.47s^{3}+1991s^{2}+4917s+6270}{s^{6}+51.21s^{5}+442.5s^{4}+2339s^{3}+6400s^{2}+9281s+6270}.$ The experimental results are also shown in Fig. 7 (b). The response of each finger is slower, and the settling time is around 1.1 $sec$. The error of each finger is larger compared to the performance of the controller with a larger bandwidth.